A company has twelve trucks, each generating $ 500 net profit per week. On average each truck will have an accident due to faulty brakes once every nine months.

These accidents result in the truck being off the road for two weeks. In addition to the repair costs averaging $ 350 each time, the trucks are not generating revenue. Damaged trucks can be replaced by hiring replacement vehicles at $ 400 per week.

Faulty brake accidents could be eliminated by more frequent maintenance checks, but this would cost an additional $ 125 per truck, per month. Eliminating these accidents would qualify the company for a reduction of 5% in its annual insurance costs of $ 2,700 per truck.

In light of the guidance to company directors and board members about leading and managing health and safety in their organisations, outline the financial case for either making this expenditure or not.
Answer #1

A Truck can have an accident a maximum of 2 times a year. To elaborate it, a truck repaired in January, February, March can have an another accident in Oct, Nov and December respectively. Information regarding last repairs of every vehicle is not available.

When no vehicle had accident in Jan, Feb and March months (then there can happen only accident in a year)

Option-1: Without Frequent Maintenance Checks and Replacement:

Profit generated by one truck in a year = 52 weeks * 500 = $26,000

Loss when a truck has an accident = Repair costs + Lost profit = 350+1000 = $1,350

Annual Insurance costs per truck per year = $2,700

Net profit per truck in a year = 26000-1350-2700= $21,950

Option-2: Without Frequent Maintenance Checks but with Replacement:

Loss when a truck has an accident = Repair costs + Lost profit = 350+400*2 = $1,150

Net profit per truck in a year = 26000-1150-2700= $22,150

Option-3: With Frequent Maintenance Checks

Additional Cost per truck due to maintenance checks monthly = $125*12 = $1,500

Annual Insurance costs per truck per year = 2700 - 5% of 2700 = $2,565

Net profit per truck in a year = 26000-1500-2565= $21,935

When all vehicle had accidents in Jan, Feb and March months

(then there are 2 accidents happening during the year per truck)

Option-1: Without Frequent Maintenance Checks and Replacement:

Profit generated by one truck in a year = 52 weeks * 500 = $26,000

Loss when a truck has an accident = Repair costs + Lost profit = (350+500*2)*2 = $1,350

Annual Insurance costs per truck per year = $2,700

Net profit per truck in a year = 26000-2700-2700= $20,600

Option-2: Without Frequent Maintenance Checks but with Replacement:

Loss when a truck has an accident = Repair costs + Lost profit = (350+400*2)*2 = $1,150

Net profit per truck in a year = 26000-2300-2700= $21,000

Option-3: With Frequent Maintenance Checks

Net profit per truck in a year (remains same)= $21,935

Suggestion: Its recommended to company's directors and board members to make the additional expenditure by way of frequent maintenance checks monthly for every truck when the historic data suggests more trucks are having accidents more than once during the year as it results in increase in net profit by $935 (21935-21000) per truck.

Its always profitable to replace the repaired trucks with hired ones.

Where the data suggests otherwise, then its suggested (financially) not to go for monthly maintenance checks since it results in decrease of net profit by $215 (22150-21935) per truck.